# EoRA: Training-free Compensation for Compressed LLM with Eigenspace Low-Rank Approximation

### v2

@@ Abstract
Abstract: In this work, we re-formulate the model compression problem into the customized compensation problem: Given a compressed model, we aim to introduce residual low-rank paths to compensate for compression errors under customized requirements from users (e.g., tasks, compression ratios), resulting in greater flexibility in adjusting overall capacity without being constrained by specific compression formats. However, naively applying SVD to derive residual paths causes suboptimal utilization of the low-rank representation capacity. Instead, we propose Training-free Eigenspace Low-Rank Approximation (EoRA), a method that directly minimizes compression-induced errors without requiring gradient-based training, achieving fast optimization in minutes using a small amount of calibration data. EoRA projects compression errors into the eigenspace of input activations, leveraging eigenvalues to effectively prioritize the reconstruction of high-importance error components. Moreover, EoRA can be seamlessly integrated with fine-tuning and quantization to further improve effectiveness and efficiency. EoRA consistently outperforms previous methods in compensating errors for compressed LLaMA2/3 models on various tasks, such as language generation, commonsense reasoning, and math reasoning tasks (e.g., $\mathbf{31.31\%}$/$\mathbf{12.88\%}$ and $\mathbf{9.69\%}$ improvements on ARC-Easy/ARC-Challenge and MathQA when compensating LLaMA3-8B that is quantized to 4-bit and pruned to 2:4 sparsity). EoRA offers a scalable, training-free solution to compensate for compression errors, making it a powerful tool to deploy LLMs in various capacity and efficiency requirements.
@@ 1 Introduction
Although Large Language Models (LLMs) exhibit superior performance across diverse applications, their empirical deployment remains challenging due to their associated considerable model size and high inference costs. To mitigate these emerging challenges, model compression research such as post-training compression (Ashkboos et al., 2024; Ma et al., 2023) and compression-aware training (Alvarez & Salzmann, 2017; Lym et al., 2019; Liu et al., 2024, 2023c) has been extensively explored to reduce the computational resource demands of serving LLMs (Zhu et al., 2023). However, most existing methods either incur significant accuracy degradation compared to uncompressed models or have high training time. Additionally, their flexibility is often limited by a discrete set of compression formats (e.g., 2:4 sparsity, 3/4-bit quantization), making it challenging to meet the diverse capacity and efficiency requirements of different users.
Figure 1: The proposed EoRA method tackles LLM model compensation by first projecting the compression error into the eigenspace of output activations. Leveraging PCA, we compute the eigenvalues, which serve as weights to prioritize columns in the weight matrix for error approximation. By adopting EoRA, users can achieve significantly better compensated compressed LLMs in just a few minutes, requiring only a small amount of calibration data. For instance, the LLaMA3-8B model pruned to $50\%$ sparsity compensated with EoRA of rank 128 can narrow the accuracy gap to the uncompressed model to as low as only $6.82\%$ and $5.2\%$ accuracy loss on ARC-C and MathQA.
To overcome the above flexibility limitation, we re-formulate the model compression problem into the customized compensation problem: Given a compressed model, we aim to introduce residual low-rank paths to compensate for compression errors under customized requirements from users, such as tasks, compression ratios, etc. Rather than focusing solely on producing compressed models with minimal performance degradation, by incorporating these residual paths, the compensated model gains greater flexibility in adjusting overall capacity, without being constrained by specific compression formats. To derive the low-rank residual paths that can represent compression errors, one straightforward method is directly decomposing compression errors with Singular Value Decomposition (SVD) (Li et al., 2024; Yao et al., 2024). However, this fails to account for the varying importance of individual model weights, resulting in suboptimal utilization of the low-rank representation capacity. Moreover, naive SVD does not guarantee the minimization of layer-wise output error (Sahu et al., 2021). Furthermore, current approaches either offer limited compensation performance by neglecting calibration data or lose flexibility due to the high computational cost of compression-aware fine-tuning, making it difficult to swiftly adjust to various tasks. This raises an important question: “How can we efficiently and effectively compensate for errors in compressed large-scale language models?
@@
To address this research question, we propose Training-free Eigenspace Low-Rank Approximation (EoRA), which retains the flexibility advantages of model compensation while enhancing both efficiency and effectiveness compared to existing approaches. To design a new error compensation framework, we first project the compression error into the eigenspace of the corresponding layer’s input activations, ensuring a direct relationship between the approximation error and the model compression loss. Inspired by the classical Principal Component Analysis (PCA) algorithm, we leverage the eigenvalues of each activation channel as importance scores to reconstruct the corresponding weight columns, as shown in Figure 1. Our method enables more effective use of the low-rank representation capacity by prioritizing the approximation of weight columns associated with larger eigenvalues while imposing lower penalties on the approximation errors of less significant columns. As a training-free optimization method, our proposed EoRA  does not require any gradient computation, achieving fast optimization in minutes using a small amount of calibration data. EoRA can also provide better initialization for fine-tuning to further enhance accuracy and offer a trade-off between accuracy and training time. Moreover, EoRA is robust to quantization which can further reduce the additional cost of residual low-rank compensation paths.
We conduct experiments on both language generation, commonsense reasoning, and math reasoning tasks to validate the effectiveness of our method for compensating the compressed LLMs. We compare our approach to the previous line of work that simply approximates the compression error with SVD on LLaMA2-7B/13B (Touvron et al., 2023a, b) and LLaMA3-8B (Dubey et al., 2024). The results demonstrate that our method significantly outperforms the plain SVD method, particularly at compensating more aggressive compressed models, achieving up to $\mathbf{3.33\%}$/$\mathbf{2.65\%}$ and $\mathbf{3.42\%}$ improvements on ARC-Easy/ARC-Challenge and MathQA when compensating 2:4 pruned LLaMA3-8B. Furthermore, we show that using EoRA for LoRA initialization consistently outperforms standard Kaiming (He et al., 2015) and SVD initialization in further fine-tuning to recover accuracy loss, narrowing the accuracy degradation for 2:4 pruned LLaMA3-8B on ARC-Easy/ARC-Challenge to only $\mathbf{4.08\%}$/$\mathbf{1.88\%}$. It can even surpass the accuracy of the original model when fine-tuning 4-bit quantized models, achieving up to $\mathbf{2.95\%}$/$\mathbf{5.04\%}$ improvement on ARC-Easy/ARC-Challenge. Additionally, to minimize the extra inference latency and memory overhead of EoRA, we show that EoRA is resilient to 3/4-bit quantization, resulting in only minimal accuracy degradation. This effectively demonstrates the practicality of using the low-rank path to compensate for compression errors.
@@
  • To overcome the flexibility limitation by conventional model compression scenarios, we re-formulate the problem into customized compensation and propose a novel method, Training-free Eigenspace Low-Rank Approximation (EoRA), which does not require any gradient computation, achieving fast optimization in minutes using a small amount of calibration data to compensate for compression errors.
  • EoRA projects the weight into the eigenspace and leverages the eigenvalues as the indicator of the importance of weight. This enables more effective use of the low-rank representation capacity than naive SVD and ensures a direct correlation between error approximation loss and layer-wise compression loss through eigenspace projection.
  • EoRA can provide better initialization for fine-tuning to further enhance accuracy, achieving or even surpassing uncompressed models. Moreover, EoRA is robust to quantization which can further reduce the additional cost of residual low-rank compensation paths.
@@ 2 Preliminaries
where $X_{l}\in\mathbb{R}^{k\times n}$ is the input activation of layer $l$ and $F$ denotes the Frobenius error between the layer-wise output. Once the compression is complete, the $W_{l}$ for each layer will be substituted with $\hat{W}_{l}$, resulting in smaller model size, faster inference, or both. However, their flexibility is often limited by a discrete set of compression formats (e.g., 2:4 sparsity, 3/4-bit quantization), making it challenging to meet the diverse capacity and efficiency requirements of different users.
To remove the constraint by specific compression formats, we re-formulate the conventional model compression problem into a customized compensation problem: Given a compressed model, we aim to introduce residual low-rank paths to compensate for compression errors under customized requirements from users, such as tasks, compression ratios, etc. With these residual paths, the compensated model gains greater flexibility in adjusting overall capacity. To derive the low-rank residual paths that can represent compression errors, one naive method is directly adopting Singular Value Decomposition (SVD) (Li et al., 2024; Yao et al., 2024). More specifically, this method relies on a closed-form solution by using SVD to approximate the compression error $\Delta W_{l}=W_{l}-\hat{W}_{l}$ as $\Delta W_{l}=U_{l}\Sigma_{l}V_{l}^{T}$, where $\Sigma_{l}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the top-$r$ largest singular value sorted in descending order, and $U_{l}\in\mathbb{R}^{d\times r}$, $V_{l}\in\mathbb{R}^{k\times r}$ are orthonormal matrices, with each column representing the singular vectors corresponding to the singular values in $\Sigma_{l}$. The product of $U_{l}$ and $\Sigma_{l}$ can then be treated as $B_{l}=U_{l}\Sigma_{l}$ with $V_{l}^{T}$ being treated as $A_{l}$. Overall, the error approximation loss can be formulated as:
$\operatorname*{arg\,min}_{B_{l},A_{l}}||\Delta W_{l}-B_{l}A_{l}||_{2}$ (2)
@@
3 Method: Training-free Eigenspace Low-Rank Approximation (EoRA)
Compared with standard model compression methods, model compensation introduces residual low-rank paths to compensate for compression errors, resulting in greater flexibility in adjusting overall capacity without being constrained by specific compression formats. However, existing methods (Li et al., 2024; Yao et al., 2024) rely mainly on plain SVD for low-rank approximation, lacking sufficient representation capacity (Barron, 1993) to fully approximate $\Delta W$. In other words, the target rank $r$ remains significantly smaller than the intrinsic rank of $\Delta W$. Therefore, it is necessary to allocate the limited representation capacity of $r$ more effectively, focusing on reconstructing the more important weights while placing less emphasis on less important segments. Moreover, naive SVD performs the approximation in the original space, failing to ensure that minimizing the approximation error (Eq. 2) directly leads to minimizing the layer-wise compression loss (Eq. 1). Furthermore, current approaches (Li et al., 2024; Yao et al., 2024) either offer limited compensation performance by neglecting calibration data or lose flexibility due to the high computational cost of compression-aware fine-tuning, making it difficult to swiftly adjust to various tasks. This raises an important question: “How can we efficiently and effectively compensate for errors in compressed LLMs?
@@ 3 Method: Training-free Eigenspace Low-Rank Approximation (EoRA)
$\operatorname*{arg\,min}_{B^{\prime},A^{\prime}}||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{2}$ (3)
and SVD is applied on $\Delta W^{\prime}$ to minimize the above equation. This loss function ensures that error columns associated with larger eigenvalues are approximated more accurately than those with smaller eigenvalues, thereby facilitating a more effective allocation of the insufficient low-rank expressive power. Since $Q$ is an orthogonal matrix, we can multiply the low-rank approximated $\Delta W^{\prime}$ with $Q^{\prime-1}=\sqrt{\Lambda}^{-1}Q^{T}$ to project back to the original space after the layer-wise reconstruction, obtaining the reconstructed error $\Delta W=\Delta W^{\prime}Q^{\prime-1}$ approximated by $B^{\prime}A^{\prime}Q^{\prime-1}$. The product of $A^{\prime}$ and $Q^{\prime-1}$ can be consolidated into a single matrix with the same dimensions as the original $A^{\prime}$, ensuring no additional inference latency as $A=A^{\prime}Q^{\prime-1}$. Then, the forward pass of the compressed model with EoRA error compensation for the input activation $X$ can be formulated as:
@@
The overall training-free optimization of Eq. 3 in EoRA can be done in minutes using only a small amount of calibration data without any gradient computation. EoRA can also provide better initialization for fine-tuning to further enhance accuracy and offer a trade-off between accuracy and training time. Moreover, EoRA is robust to quantization which can further reduce the additional cost of residual low-rank compensation paths. Please refer to Sec. 4.4 and 4.5 for more details. The overall eigenspace projection method is depicted in Figure 1 with the detailed algorithm in Alg. 1.
Algorithm 1 Training-free Eigenspace Low-Rank Approximation (EoRA)
Input: $\tilde{X}$: Average of the input activations of current layer over the calibration set, $W$: Full-precision Weight, $\hat{W}$: Compressed Weight, $r$: Compensation rank
@@ 7. The final forward pass of current layer becomes $\hat{W}X+B^{\prime}AX$
The goal of low-rank compensation is to approximate $\Delta W$. To achieve this, we reformulate the compression objective for each layer as:
@@
$\operatorname*{arg\,min}_{B,A}||\Delta WX-BAX||_{F}=\operatorname*{arg\,min}_{B,A}[\text{trace}((\Delta W-BA)XX^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}$ (6)
Directly applying SVD on $\Delta W$ initially does not guarantee the minimization of the above equation Eq.6, as dropping the smallest singular values does not necessarily lead to the smallest layer-wise compression error (Eq.6) compared to discarding other singular values. To address this issue, EoRA  projects $\Delta W$ into the eigenspace before performing SVD.
Considering the case when dropping the $i^{th}$ singular value $\sigma^{\prime}_{i}$ and its corresponding singular vectors $u^{\prime}_{i}$ and $v^{\prime}_{i}$ after projecting $\Delta W$ to the eigenspace, the corresponding layer-wise compression loss can be formulated as:
$||\Delta WX-B^{\prime}A^{\prime}Q^{\prime-1}X||_{F}=||u^{\prime}_{i}\sigma^{\prime}_{i}{v^{\prime}}_{i}^{T}Q^{\prime-1}X||_{F}$ (7)
and from Eq.6, we can then rewrite the above equation as:
$||u^{\prime}_{i}\sigma^{\prime}_{i}{v^{\prime}}_{i}^{T}Q^{\prime-1}X||_{F}=\sigma^{\prime}_{i}\cdot\text{trace}(u^{\prime}_{i}{v^{\prime}}_{i}^{T}Q^{\prime-1}XX^{T}{Q^{\prime-1}}^{T}v^{\prime}_{i}{u^{\prime}}_{i}^{T})^{\frac{1}{2}}$ (8)
Since $XX^{T}=Q^{\prime}Q^{\prime T}$ and $U^{\prime}$, and $V^{\prime}$ are orthogonal matrices, we have:
$Q^{\prime-1}XX^{T}{Q^{\prime-1}}^{T}=I;{v^{\prime}}_{i}^{T}v^{\prime}_{i}=1;\text{trace}(u^{\prime}_{i}{u^{\prime}}_{i}^{T})=1$ (9)
Then,
$||u^{\prime}_{i}\sigma^{\prime}_{i}{v^{\prime}}_{i}^{T}Q^{\prime-1}X||_{F}=\sigma^{\prime}_{i}$ (10)
This result demonstrates that truncating the $i^{th}$ singular value in the eigenspace leads to the smallest layer-wise compression error compared to discarding any other singular value. Since SVD minimizes the error approximation loss, the analysis above also reveals that eigenspace projection creates a direct connection between the error approximation loss and layer-wise model compression loss.

### v3

@@ Abstract
Abstract: In this work, we re-formulate the model compression problem into the customized compensation problem: Given a compressed model, we aim to introduce residual low-rank paths to compensate for compression errors under customized requirements from users (e.g., tasks, compression ratios), resulting in greater flexibility in balancing accuracy and overhead(inference and model size) without being bound to fixed compression formats. However, naively applying SVD to derive residual paths causes suboptimal utilization of the low-rank representation capacity. Instead, we propose Training-free Eigenspace Low-Rank Approximation (EoRA), a method that directly minimizes compression-induced errors without requiring gradient-based training, achieving fast optimization in minutes using a small amount of calibration data. EoRA projects compression errors into the eigenspace of input activations, leveraging eigenvalues to effectively prioritize the reconstruction of high-importance error components. Moreover, EoRA can be seamlessly integrated with fine-tuning and quantization to further improve effectiveness and efficiency. EoRA consistently outperforms previous methods in compensating errors for compressed LLaMA2/3 models on various tasks, such as language generation, commonsense reasoning, and math reasoning tasks (e.g., $\mathbf{31.31\%}$/$\mathbf{12.88\%}$ and $\mathbf{9.69\%}$ improvements on ARC-Easy/ARC-Challenge and MathQA when compensating LLaMA3-8B that is quantized to 4-bit and pruned to 2:4 sparsity). EoRA offers a scalable, training-free solution to compensate for compression errors, making it a powerful tool to deploy LLMs more flexibly. Code is available at https://github.com/NVlabs/EoRA.
@@ 1 Introduction
Figure 1: An overview of our proposed EoRA, which projects the compression error into the eigenspace of input activations and performs low-rank approximation for compensating the compressed model. EoRA of rank 128 can improve the accuracy of the LLaMA3-8B model pruned to $2:4$ structured sparsity and quantized to 4-bit by $31.3\%$ on ARC-E and $9.7\%$ on MathQA.
@@
Although Large Language Models (LLMs) exhibit superior performance across diverse applications, their empirical deployment remains challenging due to their associated considerable model size and high inference costs. To mitigate these emerging challenges, model compression research such as post-training compression (Ashkboos et al., 2024; Ma et al., 2023) and compression-aware training (Alvarez & Salzmann, 2017; Lym et al., 2019; Liu et al., 2024, 2023c) has been extensively explored to reduce the computational resource demands of serving LLMs (Zhu et al., 2023). However, most existing methods either incur significant accuracy degradation compared to uncompressed models or have high training time. Also, their adaptability is often constrained by a limited range of hardware-supported compression formats (e.g., 2:4 sparsity, 3/4-bit quantization), making it difficult to address various user requirements for accuracy and efficiency. For example, if a user is willing to accept slightly increased inference latency to gain better accuracy, the strict 2:4 sparsity requirement on NVIDIA GPUs or existing integer quantization kernels rules out any intermediate approach, such as 2.X:4 sparsity or INT.X-bit quantization, where X can be any arbitrary value.
To bypass such limited flexibility, we reframe model compression as the customized compensation problem: introduce residual low-rank paths to compensate for compression errors under diverse user needs (e.g., tasks, compression ratios). Instead of focusing solely on minimal performance drop, these paths boost capacity without being constrained by specific compression formats. A straightforward approach is using SVD (Li et al., 2024b; Yao et al., 2024) on compression errors, but this does not factor in the differing influence each weight matrix element has on the model’s final output, resulting in suboptimal utilization of the low-rank representation capacity. Other existing methods either yield limited compensation (by ignoring calibration data) or lose flexibility (via costly compression-aware fine-tuning), making quick task adaptation difficult. This leads to the question: “How can we efficiently and effectively compensate errors for compressed large language models?
@@
To address this research question, We introduce Training-free Eigenspace Low-Rank Approximation (EoRA), which boosts effectiveness of compensation while maintaining flexibility and efficiency. By projecting compression errors into the eigenspace of each layer’s input activations, we ensure a direct relationship between the approximation error and the model compression loss. Inspired by the classical Principal Component Analysis (PCA), our method uses each activation channel’s eigenvalues as importance scores for reconstructing weight columns, maximizing low-rank capacity. As a training-free optimization method, EoRA requires no gradient computation and can be completed within a few minutes using minimal calibration data. It can also serve as a good starting point for fine-tuning, and remains robust to quantization to further reduce overhead.
We validate the effectiveness of EoRA for compensating the compressed LLMs (LLaMA2-7B/13B and LLaMA3-8B) on language generation, commonsense reasoning, and math tasks, comparing against an SVD-based approach. Our method consistently outperforms SVD, especially for aggressively compressed (including pruned, quantized, and both) models (e.g., $\mathbf{3.33\%}$/$\mathbf{2.65\%}$ and $\mathbf{3.42\%}$ improvement on ARC-Easy/ARC-Challenge and MathQA when compensating 2:4 pruned LLaMA3-8B). To address memory access overhead from naively running low-rank compensation alongside the quantized weight, we develop a custom kernel that fuses part of the low-rank matrix multiplication with the quantization kernel, yielding up to a 1.3x speedup. Moreover, EoRA remains resilient under 3/4-bit quantization with minimal accuracy drop, underscoring its practicality in compensating for compression errors.
@@
  • Flexible Model Compression: We reframe the conventional compression problem as customized compensation problem and propose a novel method, Training-free Eigenspace Low-Rank Approximation (EoRA), a training-free approach that compensates compression errors in minutes using minimal calibration data.
  • Eigenspace Projection: EoRA projects weights into the eigenspace and uses eigenvalues as importance indicators, more effectively leveraging low-rank capacity than naive SVD while directly correlating compensation with layer-wise compression loss.
  • Efficient Inference: We develop a custom kernel that fuses part of the low-rank matrix multiplication with quantization kernel, accelerating EoRA inference by up to 1.3x. EoRA is also robust to quantization, further minimizing the memory overhead from residual low-rank compensation paths.
@@ 2 Preliminaries
where $X_{l}\in\mathbb{R}^{k\times n}$ is the input activation of layer $l$ and $F$ denotes the Frobenius error between the layer-wise output. Once the compression is complete, the $W_{l}$ for each layer will be substituted with $\hat{W}_{l}$, resulting in a smaller model size, faster inference, or both. However, their flexibility is often limited by a discrete set of compression formats (e.g., 2:4 sparsity, 3/4-bit quantization), making it challenging to meet the diverse capacity and efficiency requirements of different users.
To remove the constraint by specific compression formats, we re-formulate the conventional model compression problem into a customized compensation problem: Given a compressed model, we aim to introduce residual low-rank paths to compensate for compression errors under customized requirements from users, such as tasks, compression ratios, etc. With these residual paths, the compensated model gains greater flexibility in adjusting overall capacity. To derive the low-rank residual paths that can represent compression errors, one naive method is directly adopting Singular Value Decomposition (SVD) (Li et al., 2024b; Yao et al., 2024). More specifically, this method relies on a closed-form solution by using SVD to approximate the compression error $\Delta W_{l}=W_{l}-\hat{W}_{l}$ as $\Delta W_{l}=U_{l}\Sigma_{l}V_{l}^{T}$, where $\Sigma_{l}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the top-$r$ largest singular value sorted in descending order, and $U_{l}\in\mathbb{R}^{d\times r}$, $V_{l}\in\mathbb{R}^{k\times r}$ are orthonormal matrices, with each column representing the singular vectors corresponding to the singular values in $\Sigma_{l}$. The product of $U_{l}$ and $\Sigma_{l}$ can then be treated as $B_{l}=U_{l}\Sigma_{l}$ with $V_{l}^{T}$ being treated as $A_{l}$. Overall, the error approximation loss can be formulated as:
$\operatorname*{arg\,min}_{B_{l},A_{l}}||\Delta W_{l}-B_{l}A_{l}||_{F}$ (2)
@@
3 Method: EoRA
Compared with standard model compression methods, model compensation introduces residual low-rank paths to compensate for compression errors, resulting in greater flexibility in adjusting overall capacity without being constrained by specific compression formats. However, existing methods (Li et al., 2024b; Yao et al., 2024) rely mainly on plain SVD for low-rank approximation, lacking sufficient representation capacity (Barron, 1993) to fully approximate $\Delta W$. In other words, the target rank $r$ remains significantly smaller than the intrinsic rank of $\Delta W$. Therefore, it is necessary to allocate the limited representation capacity of $r$ more effectively, focusing on reconstructing the more important weights while placing less emphasis on less important segments. Moreover, naive SVD performs the approximation in the original space, failing to ensure that minimizing the approximation error (Eq. 2) directly leads to minimizing the layer-wise compression loss (Eq. 1). Furthermore, current approaches (Li et al., 2024b; Yao et al., 2024) either offer limited compensation performance by neglecting calibration data or lose flexibility due to the high computational cost of compression-aware fine-tuning, making it difficult to swiftly adjust to various tasks. This phenomenon raises an emerging question: “How can we efficiently and effectively compensate for errors in compressed LLMs?
@@ 3 Method: EoRA
$\operatorname*{arg\,min}_{B^{\prime},A^{\prime}}||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$ (3)
where SVD is applied on $\Delta W^{\prime}$ to minimize the above equation and $B^{\prime}$ and $A^{\prime}$ denote the corresponding solutions in the eigenspace. This loss function ensures that error columns associated with larger eigenvalues are approximated more accurately than those with smaller eigenvalues, thereby facilitating a more effective allocation of the insufficient low-rank expressive power. Since $Q$ is an orthogonal matrix, we can multiply the low-rank approximated $\Delta W^{\prime}$ with $Q^{\prime-1}=\sqrt{\Lambda}^{-1}Q^{T}$ to project back to the original space after the layer-wise reconstruction, obtaining the reconstructed error $\Delta W=\Delta W^{\prime}Q^{\prime-1}$ approximated by $B^{\prime}A^{\prime}Q^{\prime-1}$. The product of $A^{\prime}$ and $Q^{\prime-1}$ can be consolidated into a single matrix with the same dimensions as the original $A^{\prime}$, ensuring no additional inference latency as $A=A^{\prime}Q^{\prime-1}$. Then, the forward pass of the compressed model compensated with EoRA for the input activation $X$ can be formulated as:
@@
The overall training-free optimization of Eq. 3 in EoRA can be done in minutes using only a small amount of calibration data without any gradient computation. EoRA can also provide better initialization for fine-tuning to further enhance accuracy and offer a trade-off between accuracy and training time. Moreover, EoRA is robust to quantization which can further reduce the additional cost of residual low-rank compensation paths. Please refer to Sec. 4.5 and 4.7 for more details. The overall eigenspace projection method is depicted in Figure 1 with the detailed algorithm in Alg. 1.
Algorithm 1 Eigenspace low-rank approximation (EoRA)
Input: $\tilde{X}$: Average of the input activations of the current layer over the calibration set, $W$: Full-precision Weight, $\hat{W}$: Compressed Weight, $r$: Compensation rank
@@ 7. The final forward pass of current layer becomes $\hat{W}X+B^{\prime}AX$
The goal of low-rank compensation is to approximate $\Delta W$ such that the approximation also minimize Eq.1. To achieve this, we reformulate the compression objective for each layer as:
@@
$&\operatorname*{arg\,min}_{B,A}||\Delta WX-BAX||_{F}\\ &=\operatorname*{arg\,min}_{B,A}[\text{trace}((\Delta W-BA)XX^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}\end{aligned}$ (6)
Directly applying SVD on $\Delta W$ initially does not guarantee the minimization of the above equation Eq.6, as dropping the smallest singular values does not necessarily lead to the smallest layer-wise compression error (Eq.6) compared to discarding other singular values. To address this issue, EoRA  projects $\Delta W$ into the eigenspace before performing SVD. In the following, we demonstrate that minimizing Eq.3 with SVD is the same as minimizing Eq.6.
Theorem. For an activation matrix $X$, whose matrix product $XX^{T}$ has an eigendecomposition given by $XX^{T}=Q\Lambda Q^{T}$. By projecting the compression error $\Delta W$ into the eigenspace with $Q\sqrt{\Lambda}$ as $\Delta W^{\prime}=\Delta WQ\sqrt{\Lambda}$, minimizing Eq.3 via SVD becomes equivalent to minimizing Eq.6.
Proof. First, note that $XX^{T}=Q\Lambda Q^{T}$, and by substituting this into Eq.6, we get
$&[\text{trace}((\Delta W-BA)Q\Lambda Q^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}\\ &=[\text{trace}((\Delta WQ-BAQ)\Lambda(\Delta WQ-BAQ)^{T})]^{\frac{1}{2}}\end{aligned}$ (7)
Since $\Lambda=\sqrt{\Lambda}\sqrt{\Lambda}$ and $\sqrt{\Lambda}=\sqrt{\Lambda}^{T}$, the above Eq. 7 can further be rewritten as:
[…]
Let $Q^{\prime}=Q\sqrt{\Lambda}$, then Eq.8 becomes:
$&[\text{trace}((\Delta WQ^{\prime}-BAQ^{\prime})(\Delta WQ^{\prime}-BAQ^{\prime})^{T})]^{\frac{1}{2}}\\ &=[\text{trace}((\Delta W^{\prime}-BAQ^{\prime})(\Delta W^{\prime}-BAQ^{\prime})^{T})]^{\frac{1}{2}}\\ &=||\Delta W^{\prime}-BAQ^{\prime}||_{F}\end{aligned}$ (9)
where the square root of the gram matrix can be transformed back to the corresponding Frobenius norm according to (Sun, 1991). Let us set $BAQ^{\prime}=B^{\prime}A^{\prime}$, then $||\Delta W^{\prime}-BAQ^{\prime}||_{F}$ becomes $||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$. By the Eckart–Young theorem (Eckart & Young, 1936), the minimization of this Frobenious norm is achieved by running SVD on $\Delta W^{\prime}$, therefore we prove that minimizing $||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$ via SVD is equivalent to minimizing Eq.6, where low-rank approximation of $\Delta W^{\prime}$ is $\text{SVD}(\Delta W^{\prime})=B^{\prime}A^{\prime}$.